# Training Neural Networks on Data Sources with Unknown Reliability

### v2

@@ Abstract
When data is generated by multiple sources, conventional training methods update models assuming equal reliability for each source and do not consider their individual data quality during training. However, in many applications, sources have varied levels of reliability that can have negative effects on the performance of a neural network. A key issue is that often the quality of data for individual sources is not known during training. Focusing on supervised learning, this work presents a solution that aims to train neural networks on each data source for a number of steps proportional to the source’s estimated relative reliability. This way, we allow training on all sources during the warm-up, and reduce learning on less reliable sources during the final training stages, when it has been shown models overfit to noise. We show through diverse experiments, this can significantly improve model performance when trained on mixtures of reliable and unreliable data sources, and maintain performance when models are trained on reliable sources only.
@@ 1 Introduction
Data sources can have differing levels of noise but in many applications, are merged together to form a single dataset. In healthcare, for example, data sources (such as different devices or sites) are often combined together [1, 2], but may not provide the same level of data quality and could contain noisy features, incorrect labelling, or missing values. These problems, if unaddressed, can have detrimental effects on the performance and robustness of machine learning models [3, 4, 5].
In this work, we consider the supervised learning case, in which a neural network is trained on data from multiple sources that are (ideally) producing unique features and labels from the same distribution, but where some sources are producing noisy features or labels at an unknown rate [^1] Training with missing data using our method could be achieved by randomly generating features or labels in place of missing values, but this is not a focus of our work..
@@
Two possible solutions for this context naturally arise: Preprocessing can be used to remove out-of-distribution observations from training [6, 7, 8] but, this requires the user to define “out-of-distribution" for the features and labels, and eliminating data assumes nothing can be learnt from noisy examples [9]. Secondly, techniques for training neural networks on noisy data can be applied, of which many exist [10, 11]. In this case however, we are not utilising the information gained from knowing an observation’s data source – which we show is useful, especially within settings of considerable noise.
@@
We therefore propose Loss Adapted Plasticity (LAP); inspired by ideas from the literature, and loss tempering, it is a general method for training neural networks on multi-source data with mixed reliability. As we do not know the sources’ true noise level a priori, we maintain a history of the empirical risk on data from each source that is used to temper the likelihood during training. This is done such that the number of steps a model is trained on a source is proportional to its estimated reliability. Hence, a model trained with LAP will benefit from seeing examples from all sources during early training, and reduce learning on less reliable data later in training, when the model is prone to memorising noisy data points [3, 4, 12, 13]. To illustrate our method’s effectiveness, we present results on diverse settings and datasets, in which we implemented or adapted baseline code from Han et al. [14], Wang et al. [15], Li et al. [16] and Li et al. [17]. We find: (1) The proposed method, LAP, improves the performance of a neural network trained on data generated by sources with mixed reliability, and (2) maintains performance when no unreliability is present. We also show that (3) LAP is applicable to a variety of contexts and different noise types, and (4) we provide implementation details and code for further development and use (see Appendix A.1).
@@ 2 Background
Noisy labels can be introduced at any point during collection and when human experts are involved, are practically inevitable [11, 18, 19, 18]: An expert might have insufficient information [20, 21]; expert labels are incorrect [20]; the labelling is subjective [22]; or there are communication problems [23]. Further, noisy features can be introduced through data processing, or faulty measurements [16]. With each of these issues affecting data sources differently, we can observe varied data quality.
The majority of the work on noisy data focuses on noisy labels where solutions usually target a combination of four aspects [10, 11]: sample selection, model architecture, regularisation, or training loss. IDPA [15] explores instance dependent noise, where label noise depends on an observation’s features, in which they estimate the true label of confusing instances using model confidence during training. Han et al. [14] propose “Co-teaching", a sample selection method, in which two neural networks are trained simultaneously. Data for which one model achieves a low training loss is selected to “teach" the other network, as they are assumed to be more reliable. This work exploits the fact that neural networks learn clean data patterns and filter noisy instances in early training [3, 4], a fact that our work will take advantage of. Noisy inputs are additionally studied in Li et al. [16], where they present RRL, employing two contrastive losses and a noise cleaning method using model confidence during training. This approach requires modifications to the model architecture and a $k$ nearest neighbours search at each epoch. However, these methods do not consider multi-source data.
In federated learning, a server trains a global model using local updates on each source [24], but when sources are noisy, these algorithms often fail [25]. To tackle varied data quality, Li et al. [17] propose ARFL, which learns global and local weight updates simultaneously. The weighted sum of empirical risk of clients’ loss is minimised, where the weights are distributed by comparing each client’s empirical risk to the average empirical risk of the best $k$ clients. The contribution to weight updates from the clients with higher losses are minimised with respect to the updates from other clients – forcing the global model to learn more from clients that achieve lower losses.
Further, Murphy [26] discusses sensor fusion, where multiple observations from sensors with differing (but known) noise are taken, with the goal to calculate the true values. This is connected to our setting, but in our case, sources generate multiple unique observations, and source reliability is unknown.
Therefore, we are exploring a context between many; federated learning, sensor fusion, and learning with noisy data, but differing enough such that methods do not apply or under-perform compared to our source-aware solution. Nevertheless, Co-teaching, IDPA, and ARFL will form our baselines.
Motivation for new methods
@@
We observe that work has so far focused on the problem of noisy data without considering individual sources. To see why this could be helpful, consider $10$ data sources $\{s_{i}\}_{i=1}^{10}$ where one source, $s_{c}$, is producing noisy data with a probability of $0.5$ and all other sources are producing clean data. Given a new observation from the noisy source, $x^{c}\in s_{c}$, consider the probability that $x^{c}$ is noisy: $p(x^{c}\notin\mathrm{R})$, where $\mathrm{R}$ is the set of reliable data and assume we know the noise rate of $s_{c}$: $p(x\notin\mathrm{R}~{}|~{}x\in s_{c})=0.5$.
@@
When knowing the data point’s source: The probability that a data point $x^{c}$ is unreliable ($x^{c}\in s_{c}$) is $p(x^{c}\notin\mathrm{R}~{}|~{}x^{c}\in s_{c})=0.5$. Similarly, $p(x^{s\neq c}\notin\mathrm{R}~{}|~{}x^{s\neq c}\notin s_{c})=0$.
Clearly, the source value can provide information about noise that is helpful during model training.
@@ 3 Methods
Inspired by Co-teaching, ARFL, and loss tempering, we propose Loss Adapted Plasticity (LAP): Briefly, we update a source reliability score as a function of the historical training empirical risk, which is used to re-weight loss such that the more reliable a source is, the longer the model trains on data from that source. Importantly, we use the fact that when training, neural networks learn non-noisy patterns before noisy data [3, 4, 14, 27, 28, 29] and are observed fitting to noisy examples only after substantial progress is made in fitting to clean examples [27]. Therefore, in early training the empirical risk on clean data is lower than noisy data (Appendix A.2).
@@
In our setting, we have a dataset $\mathcal{D}$ collected from sources in $S$ with $\mathcal{D}=\bigcup_{s\in S}\mathcal{D}_{s}$ and $s$ as either a noisy source $s\notin\mathrm{R}$ or a non-noisy source $s\in\mathrm{R}$, where $\mathrm{R}$ represents the reliable sources. An optimal model on non-noisy data $\mathcal{D}_{R}$ (parameterised by $\theta$) would have log-likelihood: $\max_{\theta}\log p(\mathcal{D}_{R}|\theta)$. However, since $R$ is unknown during training, we propose to use the tempered log-likelihood on all data $\mathcal{D}$. With $f(\cdot):[0,+\infty)\rightarrow[1,+\infty)$ as a positive and monotonically increasing function:
$\log p(\mathcal{D}_{R}|\theta)=\sum_{s\in R}\log p(\mathcal{D}_{s}|\theta)\approx\sum_{s\in S}\frac{1}{f(C_{s})}\log p(\mathcal{D}_{s}|\theta)=\log p_{\text{temp}}(\mathcal{D}|\theta)$ (1)
Where $C_{s}$ is the number of training steps that we believe source $s$ contains noise harmful to training, minus the number of steps it is considered non-noisy (clipped at 00). Therefore during late training, $C_{s}$ is large for noisy sources and $C_{s}=0$ for non-noisy sources, giving: a large temperature $f(C_{s})>>1$ for $s\notin\mathrm{R}$, and a low temperature $f(C_{s})=1$ for $s\in\mathrm{R}$, providing an approximation of $\log p(\mathcal{D}_{R}|\theta)$.
@@
To calculate $C_{s}$, we use the assumption (discussed above) that neural networks achieve a lower empirical risk on clean data than noisy data in early training. Given a source $s$ for which we wish to update $C_{s}$, with all other sources as $s^{\prime}$; using $p_{\theta}(\cdot)=p(\cdot|\theta)$; and setting $\lambda>0$, we perform:
$\begin{array}[]{c}C_{s}=\begin{cases}C_{s}+1,&\log p_{\theta}(D_{s})<\texttt{exptvar}_{s^{\prime}}(\log p_{\text{temp},\theta}(D_{s^{\prime}}),\lambda)\\ C_{s}-1,&\text{otherwise}\end{cases}\vspace{0.5em}\\ C_{s}=\max\{0,C_{s}\}\vspace{0.5em}\\ \texttt{exptvar}_{\vartriangle}(\blacktriangle,\lambda)=\mathbb{E}_{\vartriangle}[\blacktriangle]-\lambda\sqrt{\text{Var}_{\vartriangle}[\blacktriangle]}\end{array}$ (5)
Intuitively, we increase $C_{s}$ and therefore the source’s temperature, if its log-likelihood is at most $\lambda$ standard deviations less than the expected tempered log-likelihood on all other sources. Consequently, if no sources are noisy, and the distribution of negative log-likelihoods (NLLs) from the data sources forms a normal distribution, then we incorrectly increase the temperature of a non-noisy source $s$ with probability $p_{\theta}(\hat{L_{s}}\geq\lambda)$ where $L_{s}=-\log p_{\theta}(\mathcal{D}_{s})$ is the NLL on data from source $s$, and $\hat{L_{s}}$ refers to the standardisation of the NLL using the mean and variance of the NLLs of all other sources, $L_{s^{\prime}}$. When we have noisy sources, we expect that their temperature is large, $f(C_{s})>>1$, after a number of steps that is representative of their relative noise level – naturally enabling learning from noisy sources for a number of steps that reflects their “usefulness".
Implementation details
In the following, we present the implementation of our method as gradient scaling for clarity. Given a dataset $\mathcal{D}$ of features and labels that is generated by $S$ sources, $\mathcal{S}=\{s_{1},...,s_{S}\}$. We denote a subset of $\mathcal{D}$ generated by source $s$ as $\mathcal{D}_{s}\subset\mathcal{D}$, with each data point corresponding to a single source.
On a single update step of a model: we have the perceived unreliability for each source $C=\{C_{s}\}_{1}^{S}$ (initially, all sources are considered reliable and so $C_{s}=0~{}\forall s$); a batch of features, labels, and source values from $\mathcal{D}$; and a history of training losses, $L$ with length $H$ for all sources. Here, $L\in\mathbb{R}^{S\times H}$, and $l_{H}^{s}\in L$ denotes the mean loss of data from source $s$ over the most recent batch containing data in $\mathcal{D}_{s}$ (hence, subscript $H$). A larger $C_{s}$ denotes a larger estimated noise level for source $s$. Each step, we update the value $C_{s}$ (and hence, temperature) for a source $s$ using the empirical risk of all other sources $s^{\prime}$ as in Algorithm 1. Further, we define $f(C_{s})$ from Equation 1:
$1/f(C_{s})=(1-d_{s})=1-\tanh^{2}(0.005\cdot\delta\cdot C_{s})$
This choice of $f$ has some nice properties discussed in this section and in Appendix A.3. We refer to $d_{s}$ as the depression value; and $\delta$ is the depression strength, controlling the rate of depression.
@@
During training, gradient contributions from source $s$ are then scaled as follows:
$\hat{g_{s}}=(1-d_{s})g_{s},~{}~{}d_{s}=\tanh^{2}(0.005\cdot\delta\cdot C_{s})$ (6)
@@
Where $\hat{g_{s}}$ is the gradient update contribution from a source $s$. As $(1-d_{s})$ is a scalar, this method can be interpreted as scaling gradient contributions (presented here), loss re-weighting as in Figure 1, or likelihood tempering as discussed previously in this section. Additionally, a model evidence interpretation of LAP can be found in Appendix A.4.
We use $0.005$ to scale the depression strength $\delta$ for easier interpretation, whilst the use of $\tanh^{2}$ ensures that the scaling applied to $g_{s}$ is in $(0,1]$ and small perturbations of $C_{s}$ around 00 do not have a significant effect on $\hat{g_{s}}$, making it more robust to randomness. For more detail, see Appendix A.3.

### v4

@@ Abstract
When data is generated by multiple sources, conventional training methods update models assuming equal reliability for each source and do not consider their individual data quality. However, in many applications, sources have varied levels of reliability that can have negative effects on the performance of a neural network. A key issue is that often the quality of the data for individual sources is not known during training. Previous methods for training models in the presence of noisy data do not make use of the additional information that the source label can provide. Focusing on supervised learning, we aim to train neural networks on each data source for a number of steps proportional to the source’s estimated reliability by using a dynamic re-weighting strategy motivated by likelihood tempering. This way, we allow training on all sources during the warm-up and reduce learning on less reliable sources during the final training stages, when it has been shown that models overfit to noise. We show through diverse experiments that this can significantly improve model performance when trained on mixtures of reliable and unreliable data sources, and maintain performance when models are trained on reliable sources only.
@@ 1 Introduction
Data sources can have differing levels of noise, but in many applications, they are merged to form a single dataset. For example, in healthcare, data sources can refer to different devices, sites, or human labellers, which are often combined [1, 2]. These sources may not provide the same level of data quality and could contain noisy features, incorrect labelling, or missing values. These problems, if not addressed, can have detrimental effects on the performance and robustness of machine learning models trained on the combined data [3, 4, 5].
@@
Two possible solutions for this context naturally arise: Preprocessing can be used to remove out-of-distribution observations from training [6, 7, 8] but this requires the user to define “out-of-distribution” for the features and labels, and disregarding data assumes that nothing can be learnt from these noisy examples [9]. Secondly, techniques can be applied to train neural networks on noisy data, of which many exist [10, 11]. However, these methods consider the noise in the data as a whole and do not utilise the information that can be gained from knowing the data source of an observation.
Within federated learning, work has explored the use of noisy data techniques to address differing data quality issues across data sources [12]. However, federated learning approaches are constrained by the assumption that data cannot be transferred between sources, resulting in limited performance when methods are translated to settings where data sharing is permitted.
We consider the supervised learning case, where data is produced by multiple sources that should be producing data from the same distribution, but where some sources are actually producing noisy data at an unknown rate. Our goal is to train a neural network on data in this setting that minimises error on a non-noisy test set.
@@
For this, we propose Loss Adapted Plasticity (LAP). Inspired by the current noisy data literature and likelihood tempering, LAP is a general method for training neural networks on multi-source data with mixed reliability. Since we do not know the true noise level of the sources a priori, we maintain a history of the empirical risk on data from each source that we use to temper the likelihood during training. This temperature is scheduled such that the number of steps in which a model is trained on a source is proportional to its estimated reliability. An overview of this process is given in Figure 1. Hence, a model trained with LAP will benefit from seeing examples from all sources during early training, and reduce learning on less reliable data later in training, when the model is prone to memorising noisy data points [4, 13, 14, 3]. Training a model in this way considerably reduces error on a non-noisy test set when trained on data containing noisy sources.
Figure 1: Loss Adapted Plasticity: Here, ER refers to the empirical risk.
To illustrate the effectiveness of our method, we present results in various settings and datasets, and compare our method to baselines from the literature on noisy learning [15, 16, 17, 14, 18] and federated learning [12].
Our work makes the following contributions:
  • The proposed method, LAP, improves the performance of a neural network trained on data generated from sources with mixed reliability, across a diverse set of experiments in classification, regression, and for multiple data modalities.
  • When there is no noise within a dataset, LAP maintains performance compared to the standard training method. This means that using our method does not result in less performant models when data does not contain noise.
  • We provide insights into the mechanics of LAP and how it enables improved performance for models trained with noisy data. As part of this, we present explanations of the hyperparameters introduced and offer intuitions for selecting values.
  • We provide implementation details and code [^1] To reproduce our work: https://github.com/alexcapstick/unreliable-sources; To implement our method on a new task: https://github.com/alexcapstick/loss_adapted_plasticity for further development, use in new settings, and ease of reproducibility (Appendix A.1).
2 Related work
Most of the work on noisy data is focused on noisy labels where solutions usually target a combination of four aspects [10, 11]: sample selection, model architecture, regularisation, or training loss. Han et al. [15] propose “Co-teaching”, a sample selection method in which two neural networks are trained simultaneously. Data for which one model achieves a low training loss are selected to “teach” the other network, as they are assumed to be more reliable. The authors exploit the fact that neural networks learn clean data patterns and filter noisy instances in early training [3, 4]. IDPA [16] explores instance-dependent noise, where the noise of the label depends on the characteristics of the observation, in which they estimate the true label of confusing instances using the confidence of the model during training. The authors Xia et al. [14] study how to use regularisation to prevent important weights from being updated late in training, preserving the learning of clean data patterns done in the early training stages. In contrast, the authors Wei et al. [18] modify the loss during training to study how label smoothing, and negative label smoothing, can improve model accuracy when trained with noisy labels.
Noisy inputs are additionally studied in Li et al. [17], where they present RRL, employing two contrastive losses and a noise cleaning method using model confidence during training. This approach requires modifications to the model architecture and a $k$-nearest neighbours search at each epoch.
However, these current works on noisy data do not consider that data can originate from multiple sources, which can have differing noise rates. We hypothesise that this information is important when data sources contain differing noise rates.
In federated learning, a server trains a global model using local updates on each source [19], but when sources are noisy, these algorithms often fail [20]. To address the varied data quality, Li et al. [12] propose ARFL, which learns global and local weight updates simultaneously. The weighted sum of the empirical risk of clients’ loss is minimised, where the weights are distributed by comparing each client’s empirical risk with the average empirical risk of the best $k$ clients. The contribution to weight updates from the clients with higher losses is minimised with respect to the updates from other clients – forcing the global model to learn more from clients that achieve lower losses. Here, the noise rates of individual data sources are considered during training, however the constraints of the federated learning setting limit this methods use when data can be shared.
Domain generalisation is an area of machine learning studying methods to enable trained models to generalise to unseen data distributions [21]. In this context, many different but related domains are seen during training, with the goal to produce a model that can continue to produce reliable predictions in a new test domain. To relate this field to our setting, we could consider the data sources in our context as domains, each producing the same data distribution but with different noise levels. Here, the test domain would produce clean data from the same distribution as the non-noisy data in the training domains. Since we focus on mitigating the effects of training with noisy data, when viewed through the lens of domain generalisation, our method schedules the training on domains relative to how much information they provide about the test domain. This is conceptually different from the current literature in domain generalisation but could inspire new research.
@@ 3 Background
Noisy data
Data noise can take a variety of forms, and impact both the quality of labels and features.
@@
Noisy labels can be introduced at any point during collection and, when human experts are involved, are practically inevitable [11, 22, 23]: An expert may have insufficient information [24, 25]; expert labels are incorrect [24]; the labelling is subjective [26]; or there are communication problems [27].
Further, noisy features can be introduced by poor data processing or measurement errors [17].
Noisy data and features can negatively impact the training of a neural network and its performance on a clean test set. When training a neural network, we use gradient updates to iterate towards an optimal set of model weights. This is usually done by minimising the empirical risk of model parameters $\theta$ on a dataset of samples $x,y\in\mathcal{D}$ of size $N$:
$R(\mathcal{D};\theta)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i})$
Where $R(\mathcal{D};\theta)$ is the empirical risk and $f_{\theta}$ corresponds to the neural network with parameters $\theta$. Updates to the parameters $\theta$ are guided by $\mathcal{L}(f_{\theta}(x_{i}),y_{i})$, which defines the error between a prediction $f_{\theta}(x_{i})$ and the true target $y_{i}$. When $x_{i}$ or $y_{i}$ are noisy, these parameter updates can be misleading, forcing the model to learn relationships that may not exist in the clean data. Therefore, a model trained with highly noisy data will likely not converge to the same parameters as a model trained with non-noisy data. This can result in a reduction of accuracy on non-noisy test examples and affect the performance of a model in production.
Noisy data sources
Since the mechanisms causing noisy data can affect data sources differently, we may observe varied data quality across sources.
In this setting, we have access to a dataset $\mathcal{D}$ that contains observations from multiple mutually exclusive data sources $s\in\mathcal{S}$. The dataset $\mathcal{D}$ is the union of all data from the sources $s\in\mathcal{S}$:
$\mathcal{D}=\bigcup_{s\in\mathcal{S}}\mathcal{D}_{s}$
Since each data source could be separately affected by a noisy data mechanism, this leads to a dataset $\mathcal{D}$ with a total noise rate, $\epsilon$ of:
$\epsilon=\frac{1}{N}\sum_{s\in\mathcal{S}}|\mathcal{D}_{s}|\times\epsilon_{s}$
Here, each source has its own noise rate $\epsilon_{s}\in[0,1]$ defining the proportion of the data in that source $\mathcal{D}_{s}$ that is affected by noise.
Why do we need new methods?
Currently, methods designed to mitigate the effects of noise train a neural network on noisy data to achieve minimal error on a non-noisy test dataset. We extend this by aiming to achieve minimal error on a non-noisy test dataset, after being trained on multi-source data with unknown noise levels.
@@
To see why this could be helpful, consider the $10$ data sources $\{s_{i}\}_{i=1}^{10}$ where one source, $s_{c}$, is producing noisy data with a probability of $0.5$ and all other sources are producing clean data. Given a new observation from the noisy source, $x^{c}\in s_{c}$, consider the probability that $x^{c}$ is noisy: $p(x^{c}\notin\mathrm{R})$, where $\mathrm{R}$ is the set of reliable data and assume that we know the noise rate of $s_{c}$: $p(x\notin\mathrm{R}~{}|~{}x\in s_{c})=0.5$.
@@
When knowing the source of the data point: The probability that a data point $x^{c}$ is unreliable ($x^{c}\in s_{c}$) is $p(x^{c}\notin\mathrm{R}~{}|~{}x^{c}\in s_{c})=0.5$. Similarly, $p(x^{s\neq c}\notin\mathrm{R}~{}|~{}x^{s\neq c}\notin s_{c})=0$.
Figure 2: Comparing source noise estimation with and without source knowledge.
A similar example is given in Figure 2, where we consider multiple noisy sources. Here, we demonstrate that we can gain more information on the probability of a given sample being noisy by conditioning it on the source it originates from. Therefore, we hypothesise that this additional information is useful for model training when some data sources are noisy.
@@ 4 Methods
Inspired by Co-teaching, ARFL, and loss tempering, we propose Loss Adapted Plasticity (LAP), illustrated in Figure 1. We continually calculate a reliability score for each source as a function of their historical empirical risks on the training data. This is used to re-weight their negative log-likelihood during training such that the more reliable a source is, the more influence it has when training. We take advantage of the fact that during training, neural networks learn general data patterns before noisy data and are observed to fit to noisy examples only after substantial progress is made in fitting to clean examples [3, 4, 15, 28, 29, 30]. For more details on this, see Appendix A.2.
In this section, we will use the following shorthand notation: $p_{\theta}(\cdot)=p(\cdot|\theta)$.
@@
Within the literature, tempered likelihood is often used to represent our belief about the aleatoric uncertainty of a dataset [31]. A temperature parameter, applied to the likelihood over samples from a dataset, down-weight their contributions in the calculation of the tempered posterior compared to the prior. As discussed in Kapoor et al. [31], this can lead to improved estimates of the uncertainty over the posterior parameters of a model and to improved accuracy. In our work, we use these ideas to temper each source independently. Therefore, these temperatures represent our belief in the relative uncertainty of the data from each data source. This tempered likelihood re-weights the contributions of each source to the model training.
With a dataset $\mathcal{D}$, made up of data $\mathcal{D}_{s}$ collected from sources $s\in S$, and a temperature parameter for each source $T_{s}$, the tempered log-likelihood for a model parametrised by $\theta$ is:
$\log p_{\text{temp},\theta}(\mathcal{D})=\sum_{s\in S}T_{s}\log p_{\theta}(\mathcal{D}_{s})$ (1)
However, since a priori $T_{s}$ is unknown and we do not have access to the noise levels of each source, our prior belief is that $T_{s}=1$. For this particular choice of temperatures, the tempered log-likelihood equals the non-tempered log-likelihood $\log p_{\theta}(\mathcal{D})$. Therefore, when training starts, we begin with an unmodified optimisation method.
An optimal model in our setting would achieve a maximum log-likelihood on the clean data. Since our sources contain a mix of noisy and non-noisy data points, we aim to approximate this. Defining $\mathcal{D}_{R}$ as all clean data in $\mathcal{D}$, we could set the temperature of all noisy sources to $T_{s}\rightarrow 0$, eliminating their contribution to the overall tempered likelihood. By calculating the weighted mean of the log-likelihood by source, we can arrive at the following approximation:
$\max_{\theta}\frac{1}{|S|}\log p_{\theta}(\mathcal{D}_{R})\approx\max_{\theta}\frac{1}{\sum_{s\in S}T_{s}}\sum_{s\in S}T_{s}\log p_{\theta}(\mathcal{D}_{s}\cap\mathcal{D}_{R})$ (2)
This is an approximation since some sources will contain both noisy and clean samples, and so the approximation would improve as the noise levels of the sources containing noisy examples increases and the number of non-noisy sources increases.
In practice, most sources will contain both clean and noisy examples to varying degrees. Our approach, therefore, detailed in the next section, iteratively updates $T_{s}$, allowing automatic scheduling of the temperature value.
@@
To schedule the temperature value during training, we use the historical empirical risk of the model on the batches. The more frequently a data source produces empirical risk values greater than the mean of all other sources, the smaller $T_{s}$ becomes, and the less influence that source has during training. Under the dynamics that neural networks learn clean data before noisy data, this allows us to take advantage of the approximation given in the previous section during late training, and learn from clean data within noisy sources earlier in model training.
To provide greater control over this process, we set $T_{s}=f(C_{s})$, where $f:[0,+\infty)\rightarrow(0,1]$ is a positive and monotonically decreasing function. This allows us to generalise the method to other scheduling functions and interpret the value of $C_{s}$ as a source’s perceived unreliability.
We update $C_{s}$ by considering the number of times that a given data source provides data that achieves a significantly greater negative log-likelihood (NLL) than the average of the other data sources. To do this, we maintain a history (of length $H$) of NLLs for each of the batches. When updating the value $C_{s}$ for a single source, we consider the weighted mean $\mu_{s^{\prime}}$ and standard deviation $\sigma_{s^{\prime}}$ of the NLL values of all other sources from their stored history. If the mean NLL value over the history for the source in question $\mu_{s}$ is greater than $\mu_{s^{\prime}}+\lambda\sigma_{s^{\prime}}$ for a chosen $\lambda$ (hyperparameter $\lambda>0$) then we increase $C_{s}$, otherwise we reduce it:
$\begin{array}[]{c}C_{s}=\begin{cases}C_{s}+1,&\mu_{s}>\mu_{s^{\prime}}+\lambda\sigma_{s^{\prime}}\\ C_{s}-1,&\text{otherwise}\end{cases}\end{array}$ (4)
To ensure that $C_{s}$ remains positive and within the domain of $f$, we clip its value below at 00. This means that $C_{s}$ corresponds to how much a source’s data historically deviates from those patterns learnt by the model. This translates to only decreasing the source’s temperature $C_{s}$ (and therefore its perceived unreliability), if its log-likelihood is at most $\lambda$ standard deviations less than the expected tempered log-likelihood on all other sources.
The algorithm used to calculate $C_{s}$ based on a history of losses is given in Algorithm A.4.
@@
In Appendix A.3, we show theoretically that when there is no noise in the dataset, and batches are large, our tempered likelihood training uses the same update as standard maximum likelihood training at each parameter step, with a high probability. This is more likely the larger $\lambda$ is.
This hyperparameter $\lambda$, which we refer to as the leniency, relates to the probability of incorrectly reducing the temperature of a non-noisy source when all sources are non-noisy (Appendix A.5). When a dataset contains noisy sources, we expect the temperature of those noisy sources to be small (i.e.: $f(C_{s})<<1$) after a number of training steps corresponding to their relative noise level – naturally enabling learning from noisy sources for a length of time that reflects their “usefulness”.
Implementation details
We require that our temperature, $f(C_{s})$, monotonically decreases as $C_{s}$ increases so that we reduce the impact of a source on the overall tempered likelihood as its perceived unreliability increases. We also would like $f(C_{s})$ to be robust to small changes in $C_{s}$ around 00, to ensure that fluctuations in the log-likelihood between batches do not incorrectly lead to a reduction in the temperature of a source. Because of its use as an activation function and so its wide accessibility, we therefore choose to use the following for our experiments:
$f(C_{s})=1-\tanh^{2}(0.005\cdot\delta\cdot C_{s})=1-d_{s}$
Where $\delta$, which we refer to as the depression strength, rescales the value $C_{s}$ to control the rate of reduction in temperature. In our analysis, we additionally scale $\delta$ by $0.005$ so that its value is more human-readable. We refer to $\tanh^{2}(0.005\cdot\delta\cdot C_{s})$ as the depression value $d_{s}$ for simplicity in later sections.
We find that this choice of $f$ provides some nice properties: $y=\tanh^{2}x$ has gradient of 00 at $x=0$, and is asymptotic to $y=1$, which are important qualities for calculating the model update scaling. The stationary point at $x=0$ also means that small perturbations in the unreliability of each source around $C_{s}=0$ have little effects on the scaling of model updates, which reduces the consequences of randomness in loss values from a given source and ensures that source contributions are only significantly scaled once it is clear that their loss values are consistently higher than those of the other sources. Secondly, since $y=\tanh^{2}x$ is asymptotic to $y=1$, the scale factor, $1-d_{s}$ for sources that are calculated to be unreliable eventually reduces to 00, allowing the model to “ignore” these data points as if they were masked in late training. Further, the use of $tanh$ as an activation function in other areas of machine learning mean that it is widely accessible. The value of $0.005$ used within the $\tanh^{2}$ function allows the depression strength parameter $\delta$ to be on the scale of $1$, as shown in Figure 4. This value is only used to aid interpretation of the hyperparameters and its value could easily be absorbed by $\delta$. For more detail on design decisions, see Appendix A.5.
Although we find this choice of $f(C_{s})$ to work well in practice, the LAP method is designed so that other functions can be used to update the influence of sources based on their perceived reliability. The only requirement on $f(C_{s})$ is that it is monotonically decreasing and $f:[0,+\infty)\rightarrow(0,1]$.
@@
During training, the gradient contributions from source $s$ are then scaled as follows:
$\hat{g_{s}}=f(C_{s})g_{s}=(1-d_{s})g_{s}$ (5)
@@
Where $\hat{g_{s}}$ is the gradient update contribution from a source $s$. As $(1-d_{s})$ is a scalar, this method can be interpreted as scaling gradient contributions (presented here), loss re-weighting as in Figure 3, or likelihood tempering as discussed previously in this section. Furthermore, a model evidence interpretation of LAP can be found in Appendix A.6.